OpenStudy (anonymous):

A triangle has side lengths of 12 cm, 15 cm, and 20 cm. Classify it as acute, obtuse, or right. (1 point) acute obtuse right There is not enough information.

OpenStudy (anonymous):

I wanna choose Right...

Directrix (directrix):

By the converse of the Pythagorean Theorem, if 20^2 = 15^2 + 12^2, then "right" is correct. Check the arithmetic and let us know. @zhane2015

Directrix (directrix):

What is 20*20?

OpenStudy (anonymous):

400

Directrix (directrix):

Good. What is 15 times 15?

Directrix (directrix):

You can enter "15 times 15" in the Google search box and get the product. @zhane2015

Directrix (directrix):

> I wanna choose Right... @zhane2015 Suggest you not choose until the arithmetic is done.

OpenStudy (anonymous):

225

Directrix (directrix):

Okay. I'll do 12 times 12 to get 144.

Directrix (directrix):

12 cm, 15 cm, and 20 cm. ------------------------ Question: Is 12^2 + 15^2 = 20^2 That is, is 144 + 225 = 400? Yes or No?

OpenStudy (anonymous):

no

Directrix (directrix):

Okay. Therefore, the triangle is NOT Right. That leaves Obtuse and Acute. Check out this attachment below: (comes from http://jwilson.coe.uga.edu/emt668/emat6680.folders/brooks/6690stuff/righttriangle/pythagconv.html )

OpenStudy (anonymous):

OMG you're the Best!

Directrix (directrix):

Let's test for acute. Is 400 < 225 + 144 [ Is 20^2 < 12^2 +15^2 ] Yes or No? @zhane2015

OpenStudy (anonymous):

no

OpenStudy (anonymous):

So it's obtuse

Directrix (directrix):

Correct.

OpenStudy (anonymous):

Thank You so much I really appreciate it

Directrix (directrix):

Glad to help. Learn that theorem on the chart and you'll get all this type problem correct.